OpenStudy (anonymous):

Can someone help me with this equation using long division? I will post the equation shortly.

OpenStudy (anonymous):

use a calculator

OpenStudy (anonymous):

\[(x-2)/5x+5x+2x^3+6\]

OpenStudy (anonymous):

I need to show work though...

OpenStudy (anonymous):

simplify the bottom so that you end up with (x-2) in the denominator so that you can cancel the numerator with it

OpenStudy (anonymous):

Can you draw what you did to solve it because I learn visually.

OpenStudy (anonymous):

Sorry I don't know if that's even possible you have to try it out and see if it comes out like that.

OpenStudy (anonymous):

I'm wondering if you posted the equation correctly - This doesn't look like a long division problem. Is the x-2 supposed to go into a polynomial? (It looks like (x-2/5x) + 5x + etc.

OpenStudy (anonymous):

I would expect long division for a polynomial would look like this: \[\frac{ 5x +2x^{3} +6 }{ x-1 }\] with the polynomial on top, and the simpler term on the bottom.

OpenStudy (anonymous):

I know right? I did it the other way and got the correct answer, but she said to use long the divison....here is the equation in my book - (5x+5x^2+2x^3+6)/(x-2). Maybe that might be a little more clear.

OpenStudy (anonymous):

Okay, that makes more sense to me. I don't have a stylus, so you'll have to bear with me.

OpenStudy (anonymous):

First you want to rearrange the terms of the polynomial on the bottom so they are in order from highest exponent to lowest, and set it up like a long division: x-2 goes into 2x^3 + 5x^2 + 5x + 6

OpenStudy (anonymous):

If you give me a few minutes, I'll write it out on paper and post a picture -- would that work?

OpenStudy (anonymous):

Yes thank you! :)

OpenStudy (anonymous):

Can you follow what I did?

OpenStudy (cwrw238):

theres an error on second line - should be -2*2x^2 = -4x^2

OpenStudy (anonymous):

Oh, I see what I did! Thanks so much! AHHH!

OpenStudy (cwrw238):

so instead of 3x^2 on third line it should read (5x^2 - (-4x^2)) = 9x^2

OpenStudy (anonymous):

Yes - thanks. cwrw is right. That affects the rest of the problem -- still there, lost&confused?

OpenStudy (anonymous):

Yes, I got it. Thank you guys so much! I still have a been to do, but I'll come back if I need help. :)

OpenStudy (anonymous):

Not to mention that I never subtracted -- I've been doing so many equation combining lately, and I aim for adding to eliminate the variable -- ERG!!! Ok - so I demonstrated the method, but my math is messed up.

OpenStudy (anonymous):

It's okay, I got it from here. :)

OpenStudy (anonymous):

Ok, good.

OpenStudy (anonymous):

Can you help me long divison a few other another problems?